# FedFDP: Fairness-Aware Federated Learning with Differential Privacy

### v3

@@ Abstract
Federated learning (FL) is a new machine learning paradigm to overcome the challenge of data silos and has garnered significant attention. However, federated learning faces challenges in fairness and data privacy. To address both of the above challenges simultaneously, we first propose a fairness-aware federated learning algorithm, termed FedFair. Then based on FedFair, we introduce differential privacy protection to form the FedFDP algorithm to address the trade-offs among fairness, privacy protection, and model performance. In FedFDP, we designed an fairness-aware gradient clipping technique to identify the relationship between fairness and differential privacy. Through convergence analysis, we determined the optimal fairness adjustment parameters to simultaneously achieve the best model performance and fairness. Additionally, for the extra uploaded loss values, we present an adaptive clipping method to minimize privacy budget consumption. Extensive experimental results demonstrate that FedFDP significantly outperforms state-of-the-art solutions in terms of model performance and fairness. Codes and datasets will be made public after acceptance.
1 Introduction
Federated Learning (FL) (McMahan et al. 2017) is a distributed machine learning framework that allows clients to collaboratively train a shared model without exposing their respective datasets. After training their local models, each client only transmits model parameters to the central server, rather than the original training data. The server aggregates these parameters to update the global model, which is then sent back to the clients for further training. Due to its ability to improve model performance through collaboration without the need to upload original data, overcoming data silos, FL has received significant attention in recent years (Zhu, Hong, and Zhou 2021; Lee et al. 2022; Chen et al. 2024). Particularly, many studies have achieved significant breakthroughs in terms of model performance (Wang et al. 2020; Li et al. 2020a, 2021b) and communication costs (Wang et al. 2019; Ling et al. 2024) in FL.
@@ 1 Introduction
To encourage more clients to participate in federated learning, establishing a fairness-aware federated learning algorithm is necessary (Li et al. 2021a). Current fairness-aware machine learning algorithms, such as (Dwork et al. 2012; Cotter et al. 2019), have not fully addressed the issue of client performance disparities in federated learning. As shown in Figure 1, due to non-independent and identically distributed (Non-IID) data, inconsistent client objectives (yellow squares) lead to significant performance imbalances of the collaboratively trained global model (red square) across different datasets held by various clients. An increasing number of scholars are beginning to focus on this phenomenon of unfair performance disparity (Li et al. 2020b, 2021b; Zhao and Joshi 2022).
On the other hand, recent studies have pointed out that training parameters may leak data privacy, such as recovering original data (Zhu, Liu, and Han 2019) or member inference attacks (Song, Ristenpart, and Shmatikov 2017). Even if only models or gradients are uploaded instead of original data, data leakage can still occur (Melis et al. 2022). To enhance data privacy in FL, differential privacy (DP) (Dwork, Roth et al. 2014) has become the preferred method for protecting privacy in federated learning due to its moderate computational overhead and solid mathematical foundation. Incorporating appropriate DP during the training phase can effectively prevent the accidental leakage of sensitive training information (Abadi et al. 2016; Carlini et al. 2019; Feldman 2020).
Although some works have explored fairness and differential privacy in federated learning separately, no prior research has considered them in a unified framework. For the first time, we propose a fairness-aware federated learning with differential privacy. However, balancing differential privacy protection, fairness, and model performance presents a significant challenge. Specifically, there are two main challenges: First, the clipping and noise addition steps in differential privacy can impact fairness and model utility in federated learning. Second, fair federated learning requires uploading loss information to the central server, which incurs additional privacy budget consumption.
In this paper, to address the first challenge, we analyze the relationship between the fairness loss function and differential privacy in federated learning. We design an fairness-aware gradient clipping strategy to balance fairness and differential privacy. This gradient clipping strategy not only meets the requirements for fairness but also satisfies the sensitivity control needed for differential privacy. Furthermore, through convergence analysis, we demonstrate that an optimal adjustment parameter can be found to achieve the best model performance and fairness. To address the second challenge, we design an adaptive clipping method for loss uploading, significantly reducing the privacy budget consumption in this part.
@@
  • First, we propose a fairness-aware federated algorithm named FedFair. It combines a novel loss function aimed at optimizing the objectives of fairness and model performance simultaneously.
@@
  • Further, we conducted a convergence analysis of FedFDP and found the optimal fairness parameter $\lambda$ by minimizing the upper bound of convergence. We analyzed the privacy loss of the FedFDP algorithm using the concept of Rényi Differential Privacy (RDP) to verify its compliance with differential privacy requirements.
@@
2 Related Work
Differentially private federated learning (DPFL) is widely used for privacy protection in federated learning. Many studies have focused on the trade-off between model performance, communication efficiency, and privacy budget in DPFL (Fu, Chen, and Han 2022; Zhou et al. 2023; Ling et al. 2024; Fu et al. 2024). In addition that, various fairness concepts in machine learning have been proposed, each emphasizing different aspects and addressing the needs of different stakeholders. For example, concepts like Good-Intent Fairness (Mohri, Sivek, and Suresh 2019) and Selection Fairness (Zhou, Fang, and Hui 2021) have been introduced. Further discussions on these concepts can be found in additional literature (Cong et al. 2020; Lyu et al. 2020) .
At present, the primary criterion for fairness in FL is group fairness, which aims to mitigate biases associated with sensitive attributes like gender and race. Relevant research has shown that in federated learning with differential privacy, adjusting parameters such as the clipping threshold and noise multiplier can help enhance the algorithm’s group fairness (Gu et al. 2022). There is a growing body of literature that addresses the concept of group fairness specifically within the framework of DPFL (Gálvez et al. 2021; Padala, Damle, and Gujar 2021; Gu et al. 2022, 2020; Pentyala et al. 2022).
The more predominant fairness criterion in FL is balanced performance fairness, as introduced by Li et al. (Li et al. 2020b). This approach ensures fairness by assessing the uniformity of model performance across participating clients. The concept has gained broad acceptance. Building on this, Zhao et al. (Zhao and Joshi 2022) proposed an alternative to the loss amplification mechanism, advocating for a direct weight reallocation strategy. Concurrently, Li et al. (Li et al. 2021b) concentrated on enhancing the fairness of local models within personalized FL. Our study pioneers the investigation of balanced performance fairness within the context of DPFL.
3 Preliminaries
3.1 Federated Learning and Fairness
We have defined balanced performance fairness in FL as follows.
Definition 1
(Balanced performance fairness (Li et al. 2020b)) For trained models $\mathbf{w}_{1}$ and $\mathbf{w}_{2}$, we informally recognize that model $\mathbf{w}_{1}$ provides a more fair solution to the federated learning objective in Equation (1) than model $\mathbf{w}_{2}$ if the performance of model $\mathbf{w}_{1}$ on the $N$ devices is more uniform than the performance of model $\mathbf{w}_{2}$ on the $N$ devices.
3.2 Differential Privacy
Definition 2
(Differential Privacy (Dwork, Roth et al. 2014)). $(\epsilon,\delta)$-differential privacy is achieved by a randomized mechanism $\mathcal{M}:\mathcal{X}^{n}\rightarrow\mathbb{R}^{d}$, for any two neighboring databases $D_{i},D_{i}^{\prime}\in\mathcal{X}^{n}$ that differ in only a single entry, and $\forall$ $\mathcal{S}\subseteq\text{Range}(\mathcal{R})$:
DPSGD. In gradient-based methods, DPSGD(Abadi et al. 2016) is widely used in privacy preservation. It performs per-sample gradient clipping and noise addition:
  • Firstly, computing the gradient of loss function: $g^{i,j}_{t}=\nabla F_{i}(\mathbf{w}_{t}^{i},\xi_{j})$.
  • Secondly, clipping the gradient of per-sample to get the sensitivity: $\hat{g}^{i,j}_{t}=g^{i,j}_{t}/\max\left(1,\frac{\|g^{i,j}_{t}\|}{C}\right)$, where $C$ represents the clipping bound.
  • Thirdly, adding Gaussian noise with mean 0 and standard deviation $C\sigma$ to the sum of clipped gradients, where $\sigma$ is the noise multiplier: $\tilde{g}^{i}_{t}=\frac{1}{|\mathcal{B}_{i}|}\left(\sum_{j=1}^{|\mathcal{B}_{i}|}\hat{g}^{i,j}_{t}+C\sigma\cdot\mathcal{N}(0,\mathbf{I})\right)$.
4 FedFair: Equipping FL with Fairness
$\min_{\mathbf{w}}H(\mathbf{w})=F(\mathbf{w})+\frac{\lambda}{2}\sum_{i=1}^{N}\left(F_{i}(\mathbf{w})-F(\mathbf{w})\right)^{2}.$ (4)
where $F(\mathbf{w}^{i}_{t})\triangleq\sum_{i=1}^{N}p_{i}F_{i}(\mathbf{w}^{i}_{t})$ and $F_{i}(\mathbf{w}^{i}_{t})$ is determined using the training dataset [^1] We use $F(\mathbf{w}_{i}^{t})\leftarrow F(\mathbf{w}_{t})\triangleq\sum_{i=1}^{N}p_{i}F_{i}(\mathbf{w}_{t}^{i})$, the reason for $F(\mathbf{w}_{t})\neq F(\mathbf{w}_{t},\xi_{test})$ is that the test data $\xi_{test}$ from the server cannot be used for training. from round $t-1$, as delineated in line 2 of Algorithm 2, and the $\lambda$ ($\lambda\geq 0$) is a hyperparameter used to adjust the degree of fairness  [^2] When $\lambda=0$, the algorithm becomes traditional federated learning. When $\lambda\rightarrow\infty$, the algorithm only focus on fairness. However, if $\lambda$ is too large, it will cause very steep gradients which might prevent the model from converging. Fortunately, we theoretically found an optimal value for $\lambda$. In addition, as shown in Figure 3, we experimentally confirmed the existence of the optimal $\lambda$.. Consequently, within the context of Equation (5), $F(\mathbf{w}^{i}_{t})$ is considered a constant when the gradient is being computed. So, the gradient of Equation (5) as follows:
$\nabla H_{i}(\mathbf{w}_{t}^{i})=\nabla F_{i}(\mathbf{w}_{t}^{i})+\lambda\Delta_{i}\cdot\nabla F_{i}(\mathbf{w}_{t}^{i})=(1+\lambda\Delta_{i})\nabla F_{i}(\mathbf{w}_{t}^{i}),$ (6)
where $\Delta_{i}=F_{i}(\mathbf{w}_{t}^{i})-F(\mathbf{w}_{t}^{i})$.
The stochastic gradient descent by using Equation (6) as show:
where $\eta_{i}=\eta(1+\lambda\Delta_{i})$.
Figure 2: FedFDP framework.
Therefore, our approach can be viewed as establishing an adaptive learning rate that allows the model update to achieve both performance and fairness objectives simultaneously. Based on these principles, we introduce the FedFair algorithm, as depicted in Appendix Algorithm 2.
5 FedFDP: Further Equipping FedFair with Differential Privacy
Next, we discuss the details of these two strategies.
1 Initialize $\mathbf{w}_{0},F(\mathbf{w}_{0}),C_{l}^{0}$;
2 SERVER
3       for $t\in[0,1,\cdots,T-1]$  do
4             for $i\in[1,2,\cdots,N]$ parallel  do
6
7            $\mathbf{w}_{t+1}=\sum_{i=1}^{N}p_{i}\mathbf{w}_{t+1}^{i}$;
8             $F(\mathbf{w}_{t+1})=\sum_{i=1}^{N}p_{i}\tilde{F}_{i}(\mathbf{w}_{t+1}^{i})$;
9
10      return $\mathbf{w}_{T}$;
11
12LOCAL
14       Sample randomly a batch $\mathcal{B}_{i}$ with probability $q$;
15       for $j\in[1,2,\cdots,|\mathcal{B}_{i}|]$  do
16             $\Delta_{i}^{j}=F_{i}(\mathbf{w}_{t}^{i},\xi_{j})-F(\mathbf{w}_{t})$;
17             Compute $C^{i,j}_{t}$ by Equation (11);
18             $\tilde{g}_{t}^{i,j}=C_{t}^{i,j}\cdot\nabla F_{i}(\mathbf{w}_{t}^{i},\xi_{j})$;
19
21
22      Compute clipping bound of loss $C_{l}^{i,t}$ by Equation (12) and (13) ;
23       for $j\in[1,2,\cdots,|\mathcal{B}_{i}|]$  do
24             Compute $F_{i}(\mathbf{w}_{t+1}^{i},\xi_{j})$;
25
26            $f_{i,j}=\min({C_{l}^{i,t}},\max(0,F_{i}(\mathbf{w}_{t+1}^{i},\xi_{j})))$;
27
29       return $\mathbf{w}_{t+1}^{i},\tilde{F}_{i}(\mathbf{w}_{t+1}^{i})$;
30
Algorithm 1 FedFDP
5.1 Fair-clipping Strategy for Gradient
In the general DPSGD algorithm (Abadi et al. 2016; Fu et al. 2023; Bu et al. 2022; Yang et al. 2022), per-sample clipping is an essential process for obtaining sensitivity. In order to be compatible with per-sample clipping, initially, we expand our objective function of client $i$ in FedFair, Equation (5), into Equation (8):
Next, we compute the gradient of Equation (8):
where $\Delta_{i}^{j}=F_{i}(\mathbf{w}_{t}^{i},\xi_{j})-F(\mathbf{w}_{t}^{i})$.
To ensure the algorithm is differentially private, we replace the gradient of loss function by Equation (9), and reformulate the computation of the gradient descent for client $i$ as follows:
$\mathbf{w}_{t+1}^{i}=\mathbf{w}_{t}^{i}-\frac{\eta}{|\mathcal{B}_{i}|}[\sum^{|\mathcal{B}_{i}|}_{j=1}C^{i,j}_{t}\cdot\nabla F_{i}(\mathbf{w}_{t}^{i},\xi_{j})+\sigma C\cdot\mathcal{N}(0,\mathbf{I})],$ (10)
where:
5.2 Adaptive Clipping Method for Loss
Since the calculation of $\Delta_{i}^{j}$ requires clients to upload $F_{i}(\mathbf{w}_{t}^{i})$, we need to incorporate differential privacy to ensure that the algorithm does not compromise privacy. As shown in lines 1-1 of Algorithm 1, We need to clip $F_{i}(\mathbf{w}_{t}^{i})$ to the interval $[0,C_{l}^{i,t}]$, and then add Gaussian noise with a mean of 00 and a standard deviation of $\sigma_{l}\cdot C_{l}^{i,t}$ to ensure differential privacy. However, setting an appropriate clipping bound $C_{l}^{i,t}$ for $F_{i}(\mathbf{w}_{t}^{i})$ is challenging. A large $C_{l}^{i,t}$ avoids loss of fidelity to the original values, but it also leads to large injected noise.
where
To ensure the achievement of differential privacy during the clipping process, we use the noise-added loss as the clipping bound $C_{l}^{i,t}$. Each round’s clipping bound adheres to the post-processing nature of differential privacy when $t=\{1,\dots,T\}$.
5.3 Overall Algorithm
  • DPSGD (lines 1-1): In lines 1-1, we compute the fair-clipping bound $C_{t}^{i,j}$ and the per-sample gradient $\nabla F_{i}(\mathbf{w}_{t}^{i},\xi_{j})$ to get the processed gradient $\tilde{g}_{t}^{i,j}$. Then, we add mean 0 noise to the sum of per-sample processed gradient $\tilde{g}_{t}^{i,j}$ in line 1 and execute the gradient decent.
  • DP for local loss (lines 1-1): As described in 5.2, add noise to the clipped local loss by using adaptive clipping bound $C_{l}^{i,t}$.
6 Finding Optimal Fairness Parameter
Datasets MNIST FashionMNIST CIFAR10
Accuracy Fairness Accuracy Fairness Accuracy Fairness
Table 1: FedFDP and baselines: results of test accuracy (%) and fairness ($\Psi$), while the privacy budget $\epsilon=3.52$.
6.1 Convergence Analysis
Assumption 1
Assumption 2
Assumption 3
Let $\xi_{j}$ be sampled from the $i$-th device’s local data uniformly at random. The expected squared norm of stochastic gradients is uniformly bounded, i.e., $\left\|\nabla F_{i}\left(\mathbf{w}_{t}^{i},\xi_{j}\right)\right\|\leq G$ for all $i=[1,\cdots,N]$ , $t=[1,\cdots,T-1]$ and $j=[1,\cdots,|\mathcal{B}_{i}|]$
Theorem 1
where:
  • $\hat{B}\triangleq\min_{i}|\mathcal{B}_{i}|$,
  • $\Gamma\triangleq F^{*}-\sum_{i=1}^{N}p_{i}F_{i}^{*}$,
The meaning of $\Gamma$ is consistent with (Li et al. 2020c), where a larger value of $\Gamma$ indicates that the data among different clients is more Non-IID.
proof sketch: We investigate the relationship between $\mathbb{E}\|\mathbf{w}_{t+1}-\mathbf{w}^{*}\|^{2}$ and $\mathbb{E}\|\mathbf{w}_{t}-\mathbf{w}^{*}\|^{2}$, and then use mathematical induction to obtain an upper bound for $\mathbb{E}\|\mathbf{w}_{t}-\mathbf{w}^{*}\|^{2}$. Finally, by utilizing Assumption 1, we derive Equation (14). For the detailed proof, please refer to Appendix.
Herterogeneity Scalability
Datasets MNIST FashionMNIST CIFAR10
FedAvg 93.40 |||| 1.1e11 94.17 |||| 4.1e9 85.60 |||| 3.7e9 52.61 |||| 6.1e9 51.58 |||| 2.0e9 55.90 |||| 1.5e9
SCAFFLOD 93.95 |||| 7.0e10 94.06 |||| 1.9e9 84.56 |||| 3.6e9 53.85 |||| 6.2e9 54.34 |||| 2.9e8 53.98 |||| 2.2e9
FedProx 90.50 |||| 5.3e10 94.23 |||| 3.1e9 84.38 |||| 5.6e9 51.34 |||| 4.4e9 53.67 |||| 4.5e8 54.14 |||| 2.2e9
FedDyn 91.42 |||| 5.6e10 93.08 |||| 3.8e9 85.21 |||| 9.8e10 52.14 |||| 5.5e9 50.21 |||| 1.2e10 51.86 |||| 2.0e10
ALI-DPFL 90.89 |||| 3.3e10 93.98 |||| 2.6e9 83.66 |||| 2.5e10 52.05 |||| 3.6e9 38.35 |||| 2.4e10 33.46 |||| 1.6e10
q-FFL 93.74 |||| 7.8e10 93.86 |||| 2.5e8 85.20 |||| 7.4e8 48.46 |||| 4.7e9 51.12 |||| 2.7e8 49.83 |||| 1.1e7
FedFDP 94.13 |||| 2.3e10 94.56 |||| 3.4e9 86.77 |||| 6.0e8 54.21 |||| 2.6e9 54.52 |||| 2.2e8 56.49 |||| 5.3e8
Table 2: FedFDP and baselines: results of test accuracy (%) and fairness ($\Psi$) in different heterogeneity and scalability, while the privacy budget $\epsilon=3.52$.
$\epsilon$ 1.0 2.0 3.0 4.0
Accuracy Fairness Accuracy Fairness Accuracy Fairness Accuracy Fairness
FedAvg 61.68 1.1e6 82.69 2.3e7 84.07 3.8e8 86.83 8.0e8
SCAFFLOD 62.12 1.2e6 82.36 2.3e7 83.39 2.6e9 85.41 5.6e8
FedProx 61.25 1.1e6 81.62 2.3e7 85.32 3.8e8 85.68 2.4e9
FedDyn 62.38 1.5e6 82.43 3.1e7 83.29 2.9e8 84.51 5.2e9
ALI-DPFL 61.78 1.1e6 83.01 2.3e7 84.12 3.9e8 85.58 4.5e9
q-FFL 58.99 3.6e6 80.24 9.6e7 82.92 5.5e8 85.01 3.8e9
FedFDP 63.36 1.0e6 83.15 2.1e7 85.59 9.5e7 86.97 3.8e8
[…]
6.2 Optimal Fairness Parameter $\lambda$
Based on the convergence analysis, we find the “best” fairness parameter $\lambda$ that can minimize the convergence bound. We have the following theorem to do so:
Theorem 2
There exists a computable optimal $\lambda$, denoted as $\lambda^{*}$, which ensures the fastest convergence of the Algorithm 1.
proof sketch: By reorganizing the convergence upper bound of Equation (14) into a function of $\lambda$:
By using numerical analysis methods, the minimum value $P(\lambda)$ can be solved. For the meaning of $a_{1,2,3,4,5}$ and the proof process, please refer to the Appendix.
7 Privacy Analysis
Because each client has the same privacy loss in every round. As a result, we only assess each client’s privacy loss from the perspective of a single client. Following that, we examine the privacy loss of FedPDF for the $i$-th client as follows.
Theorem 3
[…]
where $10<\delta<$ 1, $R_{model}^{i}(\alpha)$ is the RDP of $w_{t+1}^{i}$ is computed by Theorem 4 in Appendix, and $R_{loss}^{i}(\alpha$ is the RDP of $\tilde{F}_{i}(\mathbf{w}_{t+1}^{i})$ which is computed by Theorem 5 in Appendix.
The details of the proof please refer to Appendix.
8 Experiments
This section outlines our experimental framework. We detail the experimental assessment of FedFDP across dimensions of performance, fairness, heterogeneity, and scalability, including ablation experiments focusing on the fairness hyperparameter $\lambda$, privacy budget $\epsilon$, clipping bound $C$ and noise multiplier $\sigma$.
In the appendix, we also provide an experimental assessment of FedFair across the same dimensions of performance, fairness, heterogeneity, and scalability. For an introduction to baselines, default values of hyperparameters, and the running environment, please refer to the Appendix.
All tasks were repeated five times and report the mean and standard deviation.
8.1 Results
Performance and Fairness. Table 1 displays the experimental results of the FedFDP algorithm compared to baseline methods. Under a consistent privacy budget of $\epsilon=3.52$, when accuracy matches the best baseline, our method significantly enhances fairness, with improvements of 30.3%, 6.7%, and 27.8% observed across three separate datasets.
Heterogeneity. Table 2 shows that FedFDP outperforms baseline methods in accuracy while maintaining significant fairness when confronted with diverse data distributions exhibiting varying levels of heterogeneity. Except for the $\mathrm{Dir}(0.5)$ scenario, where it slightly underperforms q-FFL, FedFDP shows fairness improvements of 30.3% and 18.9% over the best baseline in $\mathrm{Dir}(0.1)$ and $\mathrm{Dir}(1.0)$, respectively.
Scalability. Table 2 demonstrates that as the number of clients increases, FedFDP consistently achieves accuracy comparable to the best baseline. Moreover, except in the 50 clients scenario where it slightly lags behind q-FFL in fairness, FedFDP exhibits fairness improvements of 27.8% and 18.5% with 10 and 20 clients, respectively.
Figure 3: Impact of different $\lambda$ in FedFDP algorithm.
Figure 4: Impact of different $C$ in FedFDP algorithm.
Figure 5: Impact of different $\sigma$ in FedFDP algorithm.
Different $\epsilon$. We conducted experiments with $\epsilon=[1,2,3,4]$ on the FashionMNIST dataset, as detailed in Table 3. For $\epsilon=[1,2,3]$, we allowed larger values of $\epsilon$ to support more iterations. At $\epsilon=4$, we supported a smaller noise multiplier $\sigma=1.65$ with a similar number of iterations as at $\epsilon=3$. The experimental results demonstrate that FedFDP consistently achieves performance on par with the best baseline in terms of accuracy across various privacy budgets. In terms of fairness, the improvements are 9.1%, 8.7%, 67.2%, and 32.1%, respectively.
9 Conclusion
Appendix A Appendix
In the beginning, Table 4 presents some notations utilized in this paper. Symbols that have not been previously introduced will be defined in subsequent sections.

### v4

@@ Abstract
Federated learning (FL) is an emerging machine learning paradigm designed to address the challenge of data silos, attracting considerable attention. However, FL encounters persistent issues related to fairness and data privacy. To tackle these challenges simultaneously, we propose a fairness-aware federated learning algorithm called FedFair. Building on FedFair, we introduce differential privacy to create the FedFDP algorithm, which addresses trade-offs among fairness, privacy protection, and model performance. In FedFDP, we developed a fairness-aware gradient clipping technique to explore the relationship between fairness and differential privacy. Through convergence analysis, we identified the optimal fairness adjustment parameters to achieve both maximum model performance and fairness. Additionally, we present an adaptive clipping method for uploaded loss values to reduce privacy budget consumption. Extensive experimental results show that FedFDP significantly surpasses state-of-the-art solutions in both model performance and fairness.
Index Terms:
Federated Learning, Differential Privacy, Fairness
I Introduction
Federated Learning (FL) [33] is a distributed machine learning framework that allows clients to collaboratively train a shared model without exposing their respective datasets. After training their local models, each client only transmits model parameters to the central server, rather than the original training data. The server aggregates these parameters to update the global model, which is then sent back to the clients for further training. Due to its ability to improve model performance through collaboration without the need to upload original data, overcoming data silos, FL has received significant attention in recent years [54, 23, 6]. Particularly, many studies have achieved significant breakthroughs in terms of model performance [46, 27, 26] and communication costs [47, 31] in FL.
@@ I Introduction
To encourage more clients to participate in federated learning, establishing a fairness-aware federated learning algorithm is necessary [25]. Current fairness-aware machine learning algorithms, such as [10, 8], have not fully addressed the issue of client performance disparities in federated learning. As shown in Figure 1, due to non-independent and identically distributed (Non-IID) data, inconsistent client objectives (orange squares) lead to significant performance disparities of the collaboratively trained global model (red square) across different datasets held by various clients. An increasing number of scholars are beginning to focus on this phenomenon of unfair performance disparity [28, 26, 50].
On the other hand, recent studies have pointed out that training parameters may leak data privacy, such as recovering original data [53] or member inference attacks [42]. Even if only models or gradients are uploaded instead of original data, data leakage can still occur [35]. To enhance data privacy in FL, differential privacy (DP) [11] has become the preferred method for protecting privacy in federated learning due to its moderate computational overhead and solid mathematical foundation. Incorporating appropriate DP during the training phase can effectively prevent the accidental leakage of sensitive training information [1, 5, 12].
Although some works have explored balanced performance fairness and differential privacy in federated learning separately, no prior research has considered them in a unified framework. For the first time, we propose a fairness-aware federated learning with differential privacy. However, balancing differential privacy protection, fairness, and model performance presents a significant challenge. Specifically, there are two main challenges: First, the clipping and noise addition processes in differential privacy have the potential to affect both fairness and model utility in federated learning. Moreover, achieving fairness in federated learning requires uploading loss information to the server, which entails additional privacy budget consumption.
In this paper, the relationship between the fairness loss function and differential privacy in federated learning is analyzed to address the first challenge. We design an fairness-aware gradient clipping strategy to balance fairness and differential privacy. This gradient clipping strategy not only meets the requirements for fairness but also satisfies the sensitivity control needed for differential privacy. Furthermore, through convergence analysis, we demonstrate that an optimal adjustment parameter can be found to achieve the best model performance and fairness. To address the second challenge, we design an adaptive clipping method for loss uploading, significantly reducing the privacy budget consumption in this part.
@@
  1. 1.
    First, a fairness-aware federated algorithm, FedFair, is proposed. It integrates a novel loss function specifically designed to simultaneously optimize fairness and model performance.
  2. 2.
@@
  3. 3.
    In addition, we conducted a convergence analysis of FedFDP and identified the fairness parameter $\lambda^{*}$ that results in the fastest convergence. We analyzed the privacy loss of the FedFDP algorithm using the concept of Rényi Differential Privacy (RDP) to verify its compliance with differential privacy requirements.
  4. 4.
@@
II Related Work
II-A Differentially Private Federated Learning
Differentially Private Federated Learning (DPFL) has become widely used to enhance privacy protection in federated learning. Enayat et al. [45] dynamically adjusted compression rates in response to training loss changes to minimize communication rounds. Zhou et al. [52] examined the impact of local iterations when using the Laplace mechanism for perturbation. Fu et al. [13] introduced adaptive noise addition, customizing it based on loss fluctuations across clients and communication rounds. Ling et al. [31] balanced communication resources and privacy budgets by identifying the locally optimal number of iterations through convergence analysis. Raouf et al. [20] improved model accuracy by updating only a fixed subset of model weights while keeping others unchanged.
II-B Fairness in DPFL
The above work in DPFL mainly focuses on model performance and privacy budget efficiency, without addressing fairness issues. In machine learning, a range of fairness concepts has been proposed, each emphasizing different aspects to meet diverse stakeholder needs. For instance, Good-Intent Fairness [38] and Selection Fairness [51] have been proposed. For a deeper understanding of these concepts, please refer to the relevant literature [7, 32, 34, 41].
In contrast, some recent studies have begun exploring group fairness within DPFL, aiming to reduce biases related to sensitive attributes such as gender and race [15, 39, 16, 17, 40]. Research suggests that adjusting parameters like clipping thresholds and noise multipliers can enhance group fairness in DPFL algorithms [16]. Under the framework of DP, adjacent datasets are defined as datasets $D$ and $D^{\prime}$ that differ by a single element, ensuring that an adversary cannot determine whether the output originates from $D$ or $D^{\prime}$. This mechanism significantly enhances the protection of attribute privacy and effectively reduces potential discrimination risks. Jagielski et al. [18] pioneered research on fair learning under DP constraints, addressing the need for predictive models to avoid discrimination based on specific attributes. They designed two learning algorithms that ensure both DP and equal opportunity, addressing challenges related to the collection of sensitive attributes. Tran et al. [44] utilized DP and Lagrangian duality to design neural networks capable of addressing fairness constraints while protecting the privacy of sensitive attributes. This approach ensures compliance with legal and ethical neutrality requirements when handling data related to gender, race, or age. Similarly, Ding et al. [9] proposed a classification model that combines a functional mechanism with decision boundary fairness to achieve both fairness and DP. This approach addresses privacy concerns related to training data and mitigates unfair practices stemming from attributes such as gender and race.
II-C Balanced Performance Fairness
Unlike the studies mentioned above, our research shifts focus to a relatively unexplored area within DPFL—balanced performance fairness. While group fairness emphasizes reducing performance biases related to sensitive attributes across different data types and is primarily used in centralized machine learning, balanced performance fairness, introduced by Li et al. [28], is more suited to federated learning. It emphasizes balancing the performance of global model across all clients, which is especially meaningful for data providers like companies and institutions. Building on the work of [28], Zhao et al. [50] proposed a strategy that replaces loss amplification with direct weight reallocation. Meanwhile, Li et al. [26] concentrated on improving fairness in personalized federated learning for local models.
Incorporating DP into model updates, with parameter adjustments such as clipping bounds and noise multipliers, can influence group fairness to some extent [16]. However, achieving balanced performance fairness necessitates adjustments to the direction of model updates. Therefore, determining how to update the model under DP constraints to ensure balanced performance fairness, while accounting for the effects of clipping and noise addition, presents a significant challenge. To the best of our knowledge, this is the first study to address balanced performance fairness within the context of DPFL.
III Preliminaries
In the beginning, Table I presents some notations utilized in this paper. Symbols that have not been previously introduced will be defined in subsequent sections.
TABLE I: Summary of main notations